OpenStudy (anonymous):

4x + 2(3x – 9) = x A. x = –2 B. x = 1.8 C. x = 2 D. x = 18

OpenStudy (kidrah69):

Step 1: Simplify both sides of the equation. 4x+2(3x−9)=x Simplify: (Show steps) 10x−18=x Step 2: Subtract x from both sides. 10x−18−x=x−x 9x−18=0 Step 3: Add 18 to both sides. 9x−18+18=0+18 9x=18 Now you just divide to get the answer

OpenStudy (anonymous):

D?

OpenStudy (kidrah69):

What is 18/9 ???

OpenStudy (anonymous):

9

jagr2713 (jagr2713):

what is \[18\div 9\]

OpenStudy (kidrah69):

NO

OpenStudy (kidrah69):

9 x ? = 18 ?

OpenStudy (anonymous):

2

OpenStudy (kidrah69):

ep :)

OpenStudy (kidrah69):

*yep

OpenStudy (kidrah69):

Good job!

jagr2713 (jagr2713):